OpenStudy (anonymous):

Given the linear equation 3x + y = 5, find the slope of its graph. A. 5 B. 3 C. –5 D. –3

OpenStudy (anonymous):

Give me the equation so that next time i know what to do. This is a practice sheet that i'm turning in for a grade. Please help!

OpenStudy (anonymous):

put them in to y =mx+b form m is the slope

OpenStudy (anonymous):

okay could you help me work out the problem?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks. now where do i start?

OpenStudy (anonymous):

so you have 3x +y = 5 to put it in to y = mx+b, all you need to do is subtract 3x to both sides

OpenStudy (anonymous):